two cards are drawn from a well shuffled ordinary deck of 52 cards .find the probability that they are both aces. if the first card is (i) replaced (ii) not replaced

Answer by Boreal(15235)

replacement: 4/52*4/52 or 1/169
not replaced 4/52*3/51=1/13*1/17=1/221